OpenStudy (jenniferjuice):

What is the slope-intercept form equation of the line that passes through (1, 3) and (3, 7)? y = −2x + 1 y = −2x − 1 y = 2x + 1 y = 2x − 1

OpenStudy (mathmate):

Assume an equation of the form y=ax+b. Calculate the slope \[a=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (jenniferjuice):

a = 4/2

OpenStudy (jenniferjuice):

what is b?

OpenStudy (mathmate):

That's good... a=4/2=2. Next, put (x1,y1) into the (x,y) of the assumed equation and solve for b

OpenStudy (jenniferjuice):

wait what

OpenStudy (mathmate):

We assumed the equation to be y=ax+b.

OpenStudy (jenniferjuice):

how to i find the equation of b?

OpenStudy (mathmate):

so put (1,3) into the equation, i.e. put x=1, y=3. Only b will be left as an unknown. So solve for b.

OpenStudy (jenniferjuice):

b=3 ?

OpenStudy (mathmate):

(1,3) in y=2x+b (a=2) x=1, y=3, so 3=2(1)+b what is b?

OpenStudy (jenniferjuice):

3-b=2 ?

OpenStudy (jenniferjuice):

b = 1 ?

OpenStudy (mathmate):

So what is b? You need the value of b to find which choice to make!

OpenStudy (jenniferjuice):

b=-1

OpenStudy (mathmate):

3=2(1)+b 3=2+b Subtract 2 from each side, 3-2=2-2+b What would the value of b?

OpenStudy (jenniferjuice):

-1!

OpenStudy (mathmate):

The next line is 3-2=2-2+b 1=0+b 1=b